# Mirage: Evaluating and Explaining Inductive Reasoning Process in Language Models

### v1

@@ Abstract
Inductive reasoning is an essential capability for large language models (LLMs) to achieve higher intelligence, which requires the model to generalize rules from observed facts and then apply them to unseen examples. We present Mirage, a synthetic dataset that addresses the limitations of previous work, specifically the lack of comprehensive evaluation and flexible test data. In it, we evaluate LLMs’ capabilities in both the inductive and deductive stages, allowing for flexible variation in input distribution, task scenario, and task difficulty to analyze the factors influencing LLMs’ inductive reasoning. Based on these multi-faceted evaluations, we demonstrate that the LLM is a poor rule-based reasoner. In many cases, when conducting inductive reasoning, they do not rely on a correct rule to answer the unseen case. From the perspectives of different prompting methods, observation numbers, and task forms, models tend to consistently conduct correct deduction without correct inductive rules. Besides, we find that LLMs are good neighbor-based reasoners. In the inductive reasoning process, the model tends to focus on observed facts that are close to the current test example in feature space. By leveraging these similar examples, the model maintains strong inductive capabilities within a localized region, significantly improving its deductive performance.
@@ 1 Introduction
Inductive reasoning, known as the ability of an intelligent agent to infer abstract rules from limited observations and apply them to new examples, is crucial for large language model (LLMs) progressing toward artificial general intelligence (AGI) (Xu et al., 2024b; Sun et al., 2024; Wang et al., 2024b). As illustrated in Figure 1, given a set of observed facts, inductive reasoning process expect the model to generate abstract rules from the provided facts (i.e. [A,B,C] $\rightarrow$ [B+C,B+C,C] in the inductive task) and apply these rules to answer specific new questions (i.e. [3,4,7] $\rightarrow$ [11,11,7] in the deductive task). Despite its significant research value, it has been relatively neglected compared to other types of reasoning (e.g., math reasoning, multi-hop reasoning, etc.).
@@
Recently, some works have started to explore this problem. They primarily evaluate the model’s inductive reasoning capabilities using various datasets (Shao et al., 2024; Cheng et al., 2024; Qiu et al., 2024; Jiang et al., 2024). Though they have made great progress, their works still have two main limitations: (1) Previous works lack comprehensive evaluation. Most works have only one evaluation task: the inductive task on collected rules (Yang et al., 2024b; Shao et al., 2024) or the deductive task on specific test samples (Chollet, 2019; Xu et al., 2024a; Qiu et al., 2024). Therefore, they can only evaluate the rule induction performance or final results of inductive reasoning, instead of comprehensively analyzing the whole process (i.e. inductive + deductive). (2) Previous works lack flexible test data. Most former datasets evaluate the overall performance of models by collecting observation and test examples under the same rules (Rule, 2020; Kim et al., 2022; Lake et al., 2019). However, due to the absence of transformation rules, it is impossible to extend these examples, resulting in a fixed test set. This limitation makes it challenging to assess the impact of factors such as distribution, quantity, and form of input examples on the model’s inductive reasoning, thereby hindering a deeper analysis of the model’s reasoning mechanisms.
@@
Based on our dataset, we perform a deeper analysis of the model’s inductive reasoning process, from which we draw two new conclusions about the inductive reasoning mechanisms of LLMs: (1) Language models are poor rule-based reasoners. As shown in the left column of Figure 1, in the rule-based reasoning paradigm, inductive reasoning involves first deriving the correct rule through the observation of examples and then using the inductive rule to answer new questions (like what humans do). However, we find that LLMs perform poorly in this paradigm: In many cases, though they can not induce a correct rule, they can still perform well on deductive tasks. Through experimentation, we observe this performance gap between induction and deduction across different prompting methods, models, observed example numbers, and scenarios. This indicates that the final performance of the LLM’s inductive reasoning rarely relies on the intermediate inductive rules. (2) Language models are good neighbor-based reasoners. Furthermore, we identify an important mechanism behind LLM’s inductive reasoning, which we refer to as “neighbor-based reasoning”: If some observed facts are close to the test examples in feature space, the model tends to leverage this similarity to improve performance on deductive tasks. For example, as shown in the right column of Figure 1, even when the model cannot generate the correct rule, it can rely on the neighbor fact [3,3,7] $\rightarrow$ [10,10,7] (here the distance between [3,3,7] and [3,4,7] is small, so we refer to them as neighbors) to successfully performs deductive tasks. We demonstrate that this paradigm persists across different scenarios, models, and observed example numbers. However, it can only enhance the performance within a localized scope.
To sum up, the main contributions of our work are as follows: (1) We present a new dataset Mirage, through it, we can comprehensively evaluate the LLM’s inductive reasoning process under more flexible settings. (2) We prove that LLM is a poor rule-based inductive reasoner. In many cases, it does not rely on inductive rules to make correct deductions. (3) We prove that LLM is a neighbor-based inductive reasoner. When performing inductive reasoning, models rely on the neighbor facts in the observed fact set to get better performance.
@@ 2.1 Rule Generation
where ${\mathbb{X}}$ is the observed fact set under the rule $f$. We believe that $f$ is the core of the problem, as it allows us to generate additional facts for ${\mathbb{X}}$ based on the rule automatically. Conversely, inferring $f$ from ${\mathbb{X}}$ requires significantly more effort due to the vast range of possible rules. Therefore, we first consider automating these rules’ large-scale synthesis.
Based on previous representative datasets (Chollet, 2019; Rule, 2020; Xu et al., 2024a), we summarize the main types of rules, resulting in five atomic operations in this dataset:
  • Add: The operation adds certain components together. For example: $[x,y,z]\rightarrow[x,x+y,z]$.
  • Copy: The operation copies some components to others. For example: $[x,y,z]\rightarrow[x,x,z]$.
  • Map: The operation applies a linear transformation to some components. For example: $[x,y,z]\rightarrow[x,ky+b,z]$. Here, to avoid the interference of complex mathematical calculations, we have $k\in[1,9]$ and $b\in[0,9]$.
  • Pad: The operation fills certain components with constant values. For example: $[x,y,z]\rightarrow[x,c,c]$, where $c\in[0,9]$.
  • Swap: The operation swaps certain components. For example: $[x,y,z]\rightarrow[z,y,x]$.
@@ 2.4 Question Generation
So far, we have constructed all the metadata that we need to generate specific questions. It is worth noting that both ${\mathbb{F}}$ and ${\mathbb{X}}$ contain only abstract rules and facts, without any specific context. Therefore, they represent the fundamental inductive reasoning test data, which is why we refer to them as meta-rules and meta-facts. As shown in Figure 2, to evaluate the practical inductive reasoning capability of models, we apply these metadata to various scenarios to generate concrete problems. Specifically, we have:
  • List Transformation (LT): List transformation is the primary format used in previous inductive reasoning tasks (Rule, 2020; Xu et al., 2024a; Chollet, 2019), and here we adopt this approach as well. We transform all fact vectors into one-dimensional lists and require the model to inductively infer the transformation rules applied to these lists.
  • Real-world Problem (RP): Previous datasets lack tests for inductive reasoning capabilities in real-world scenarios (Rule, 2020; Xu et al., 2024a; Qiu et al., 2024). [^2] Here, real-world scenarios refer to mathematical inductive reasoning within natural language contexts. To mitigate this gap, we populate the metadata into different natural language templates across five real-life scenarios. The example in Figure 2 describes a trading scenario, where we use different items to represent different dimensions of the vector. All item transactions follow the same rule.
  • Code Generation (CG): For each fact, we use ${\bm{x}}$ as the input and ${\bm{y}}$ as the output of a function. The model is then tasked with predicting the corresponding Python function.
  • String Transformation (ST): The former three scenarios are related to numbers. Here, we replace the basic elements in the fact vectors with characters to conduct a new test. Notably, we modify the operations as follows: addition in the Add and Map operations is replaced with string concatenation, multiplication in Map is replaced with character replication, zero-padding in Pad becomes character deletion, and the numbers 00-$9$ are replaced with the characters $a$-$j$.
For each scenario, we design both inductive (Ind) and deductive (Ded) tasks. In the inductive task, given a set of observed facts ${\mathbb{X}}$, we ask the model to generate the rule $f$ in a specified format (e.g. a Python Function for CG) and then evaluate the accuracy of the generation. In the deductive task, using the same observed facts, we provide an unseen example input ${\bm{x}}_{t}$ as input and measure the accuracy of the predicted ${\bm{y}}_{t}$. We provide all the prompts used for these tasks in Appendix A.3.
@@ 3.1 Overall Performances on Mirage
Llama2-13B Ind 0.01 0.00 0.00 0.03 0.01 0.01 0.00 0.21 0.00 0.01 0.00 0.10
Ded 0.26 0.11 0.25 0.22 0.13 0.03 0.14 0.25 0.06 0.01 0.06 0.19
Llama3-8B Ind 0.15 0.11 0.19 0.19 0.23 0.04 0.14 0.22 0.16 0.02 0.08 0.21
Ded 0.30 0.15 0.25 0.25 0.20 0.12 0.25 0.29 0.09 0.11 0.16 0.24
GPT-4o Ind 0.41 0.32 0.38 0.32 0.35 0.21 0.44 0.30 0.33 0.16 0.41 0.24
Ded 0.68 0.37 0.61 0.56 0.58 0.25 0.64 0.39 0.42 0.17 0.49 0.29
GPT-4 Ind 0.47 0.29 0.41 0.28 0.58 0.22 0.56 0.27 0.46 0.15 0.45 0.23
Ded 0.68 0.37 0.61 0.57 0.63 0.29 0.71 0.44 0.42 0.21 0.64 0.30
Claude-3.5 Ind 0.44 0.35 0.34 0.46 0.22 0.20 0.38 0.33 0.24 0.13 0.38 0.26
Ded 0.79 0.45 0.62 0.58 0.65 0.33 0.76 0.45 0.46 0.24 0.59 0.30
Table 1: Overall performance of different models on Mirage. The best results in the inductive task (Ind) are in bold, while the best results in the deductive task (Ded) are underlined.
@@ Results
The results are shown in Table 1, from which we can draw the following conclusions: (1) LLMs’ deductive reasoning does not rely on rule induction. Given the same set of observed facts, the model’s performance on inductive tasks is noticeably worse than on deductive tasks in almost all cases. This suggests that most of the model’s correct deductions do not depend on inducing a correct rule. (2) LLMs face difficulties in handling inductive reasoning in real-world problems. When comparing different scenarios, all models perform the worst on the RP tasks. For example, GPT-4o only achieves 0.16 and 0.17 accuracy when the dimension is 8. This indicates that, compared to purely symbolic forms (LT, CG, ST), natural language forms pose a greater challenge for the models’ inductive reasoning abilities.
@@
In the main experiments, we find that there is a significant performance gap between inductive and deductive tasks for LLMs. However, this gap may be caused by the difference in difficulty between the two tasks. When the model is unable to perform correct inductive reasoning,
Model Inductive Deductive
@@
it is likely to guess the correct answers for the deductive task more easily compared to the inductive task, resulting in a higher accuracy. We conduct this additional experiment to eliminate the interference of this factor. Specifically, we randomly perturb one fact in ${\mathbb{X}}$ to violate rule $f$. Then, we observe the performance of both tasks and calculate the change rate (CR) of accuracy before and after the perturbation. CR represents the sensitivity of the model’s performance to the input. If CR is high, it indicates a strong correlation between task performance and input, making it difficult to answer the question correctly through random guessing, Therefore, CR can serve as an indicator of the reasoning difficulty for the task. We randomly choose 100 pieces of test data from the dataset and generate questions under the LF scenario. [^4] Unless otherwise specified, this configuration will be maintained for all subsequent experiments. The experimental results on different models are demonstrated in Table 2. We can observe that the two tasks have comparable CR for both models, indicating that the reasoning difficulty of the deductive task is not lower than the inductive task. The tasks themselves do not cause such a large performance gap.
@@ 3.2 Performances of Advanced Methods
In §$\S$3.1, we observe that LLMs perform poorly on our dataset, especially in inductive tasks. Considering previous work has proposed numerous methods to elicit the model’s reasoning abilities (Wei et al., 2022; Wang et al., 2023b; Madaan et al., 2023), we wonder whether they can boost models’ performance on Mirage.
@@
Since we focus on exploring the model’s intrinsic capabilities, we only consider methods that do not introduce any external tools or knowledge. Specifically, the methods are as follows: Input-Output (IO): We prompt models to generate answers directly under different shots. Inductive-Deductive (ID): We prompt models to generate rules for inductive tasks and apply them to answer questions in deductive tasks. Chain-of-Thought (CoT) (Wei et al., 2022): We prompt models to generate rationales and answers for the two tasks. Self-Consistency (SC) (Wang et al., 2023b): Based on CoT, we sample $n$ rationales and use the major voting strategy to predict the final answer. Self-Refine (SR) (Madaan et al., 2023): We prompt the model to provide feedback on the generated rules, and then refine the rules based on that feedback (with a maximum of $t$ iterations). After the iteration stops, we use the latest rule to answer the inductive task and apply it to answer the deductive task. Hypothesis Refinement (HR) (Qiu et al., 2024): HR is an optimized version of SR, which first generates $n$ rules. In each iteration, we apply the current rules to all observed examples, compare the actual output with the expected output, and get the number of correct predictions along with the information about incorrect examples. If a candidate rule is correct for all observed facts, it is immediately returned. Otherwise, the rule with the highest number of correct predictions and the associated error information is used as input for the model to refine, generating $n$ rules for the next round, until the maximum number of iterations $t$ is reached. We sample 200 questions for each test.
@@
Ind Ded ($\Delta$) Ind Ded ($\Delta$) Ind Ded ($\Delta$) Ind Ded ($\Delta$)

### v2

@@ Abstract
Inductive reasoning is an essential capability for large language models (LLMs) to achieve higher intelligence, which requires the model to generalize rules from observed facts and then apply them to unseen examples. We present Mirage, a synthetic dataset that addresses the limitations of previous work, specifically the lack of comprehensive evaluation and flexible test data. In it, we evaluate LLMs’ capabilities in both the inductive and deductive stages, allowing for flexible variation in input distribution, task scenario, and task difficulty to analyze the factors influencing LLMs’ inductive reasoning. Based on these multi-faceted evaluations, we demonstrate that the LLM is a poor rule-based reasoner. In many cases, when conducting inductive reasoning, they do not rely on a correct rule to answer the unseen case. From the perspectives of different prompting methods, observation numbers, and task forms, models tend to conduct correct deduction without correct inductive rules consistently. Besides, we find that LLMs are good neighbor-based reasoners. In the inductive reasoning process, the model tends to focus on observed facts that are close to the current test example in feature space. By leveraging these similar examples, the model maintains strong inductive capabilities within a localized region, significantly improving its reasoning performance.
@@ 1 Introduction
Inductive reasoning, known as the ability of an intelligent agent to infer abstract rules from limited observations and apply them to new examples, is crucial for large language model (LLMs) (Li et al., 2023; Wen et al., 2023; Sun et al., 2024a) progressing toward artificial general intelligence (AGI) (Xu et al., 2024b; Sun et al., 2024b; Wang et al., 2024b). As illustrated in Figure 1, given a set of observed facts, inductive reasoning process expect the model to generate abstract rules from the provided facts (i.e. [A,B,C] $\rightarrow$ [B+C,B+C,C] in the rule induction task) and apply these rules to answer specific new questions (i.e. [3,4,7] $\rightarrow$ [11,11,7] in the example inference task). Despite its significant research value, it has been relatively neglected compared to other types of reasoning (e.g., math reasoning, multi-hop reasoning, etc.).
@@
Recently, some works have started to explore this problem. They primarily evaluate the model’s inductive reasoning capabilities using various datasets (Shao et al., 2024; Cheng et al., 2024; Qiu et al., 2024; Jiang et al., 2024). Though they have made great progress, their works still have two main limitations: (1) Previous works lack comprehensive evaluation. Most works have only one evaluation task: the inductive task on collected rules (Yang et al., 2024b; Shao et al., 2024; Wang et al., 2025) or the deductive task on specific test samples (Chollet, 2019; Xu et al., 2024a; Qiu et al., 2024). Therefore, they can only evaluate the rule induction performance or final results of inductive reasoning, instead of comprehensively analyzing the whole process (i.e. inductive + deductive). (2) Previous works lack flexible test data. Most former datasets evaluate the overall performance of models by collecting observation and test examples under the same rules (Rule, 2020; Kim et al., 2022; Lake et al., 2019). However, due to the absence of transformation rules, it is impossible to extend these examples, resulting in a fixed test set. This limitation makes it challenging to assess the impact of factors such as distribution, quantity, and form of input examples on the model’s inductive reasoning, thereby hindering a deeper analysis of the model’s reasoning mechanisms.
@@
Based on our dataset, we perform a deeper analysis of the model’s inductive reasoning process, from which we draw two new conclusions about the inductive reasoning mechanisms of LLMs: (1) Language models are poor rule-based reasoners. As shown in the left column of Figure 1, in the rule-based reasoning paradigm, inductive reasoning involves first deriving the correct rule through the observation of examples and then using the inductive rule to answer new questions (like what humans do). However, we find that LLMs perform poorly in this paradigm: In many cases, though they can not induce a correct rule, they can still perform well on example inference tasks. Through experimentation, we observe this performance gap between induction and deduction across different prompting methods, models, observed example numbers, and scenarios. This indicates that the final performance of the LLM’s inductive reasoning rarely relies on the intermediate inductive rules. (2) Language models are good neighbor-based reasoners. Furthermore, we identify an important mechanism behind LLM’s inductive reasoning, which we refer to as “neighbor-based reasoning”: If some observed facts are close to the test examples in feature space, the model tends to leverage this similarity to improve inductive reasoning performance. For example, as shown in the right column of Figure 1, even when the model cannot generate the correct rule, it can rely on the neighbor fact [3,3,7] $\rightarrow$ [10,10,7] (here the distance between [3,3,7] and [3,4,7] is small, so we refer to them as neighbors) to successfully performs the reasoning. We demonstrate that this paradigm persists across different scenarios, models, and observed example numbers. However, it can only enhance the performance within a localized scope.
To sum up, the main contributions of our work are as follows: (1) We present a new dataset Mirage, through it, we can comprehensively evaluate the LLM’s inductive reasoning process under more flexible settings. (2) We find that LLM is a poor rule-based inductive reasoner. In many cases, it does not rely on inductive rules to make correct deductions. (3) We demonstrate that LLM is a neighbor-based inductive reasoner. When performing inductive reasoning, models rely on the neighbor facts in the observed fact set to get better performance. Our code is available at: https://github.com/BugMakerzzz/mirage.
@@ 2.1 Rule Generation
where ${\mathbb{X}}$ is the observed fact set under the rule $f$. Here $f$ is the core of the problem, as it allows us to generate facts for ${\mathbb{X}}$ based on it automatically. Conversely, inferring $f$ from ${\mathbb{X}}$ requires significantly more effort due to the vast range of possible rules. Therefore, we first consider automating these rules’ large-scale synthesis.
Based on previous representative datasets (Chollet, 2019; Rule, 2020; Xu et al., 2024a), we summarize the main types of rules, resulting in five atomic operations in this dataset: (1) Add: The operation adds certain components together. For example: $[x,y,z]\rightarrow[x,x+y,z]$. (2) Copy: The operation copies some components to others. For example: $[x,y,z]\rightarrow[x,x,z]$. (3) Map: The operation applies a linear transformation to some components. For example: $[x,y,z]\rightarrow[x,ky+b,z]$. To avoid the interference of complex math calculations, we have $k\in[1,9]$ and $b\in[0,9]$. (4) Pad: The operation fills certain components with constant values. For example: $[x,y,z]\rightarrow[x,c,c]$, where $c\in[0,9]$. (5) Swap: The operation swaps certain components. For example: $[x,y,z]\rightarrow[z,y,x]$.
@@ 2.4 Question Generation
So far, we have constructed all the metadata that we need to generate specific questions. It is worth noting that both ${\mathbb{F}}$ and ${\mathbb{X}}$ contain only abstract rules and facts, without any specific context. Therefore, they represent the fundamental inductive reasoning test data, which is why we refer to them as meta-rules and meta-facts. As shown in Figure 2, to evaluate the practical inductive reasoning capability of models, we apply these metadata to various scenarios to generate concrete problems. Specifically, we have: (1) List Transformation (LT): List transformation is the primary format used in previous inductive reasoning tasks (Rule, 2020; Xu et al., 2024a; Chollet, 2019), and here we adopt this approach as well. We transform all fact vectors into one-dimensional lists and require the model to inductively infer the transformation rules applied to these lists. (2) Real-world Problem (RP): Previous datasets lack tests for inductive reasoning capabilities in real-world scenarios (Rule, 2020; Xu et al., 2024a; Qiu et al., 2024). [^2] Here, real-world scenarios refer to mathematical inductive reasoning within natural language contexts. To mitigate this gap, we populate the metadata into different natural language templates across five real-life scenarios. The example in Figure 2 describes a trading scenario, where we use different items to represent different dimensions of the vector. All item transactions follow the same rule. (3) Code Generation (CG): For each fact, we use ${\bm{x}}$ as the input and ${\bm{y}}$ as the output of a function. The model is then tasked with predicting the corresponding Python function. (4) String Transformation (ST): The former three scenarios are related to numbers. Here, we replace the basic elements in the fact vectors with characters to conduct a new test. Notably, we modify the operations as follows: addition in the Add and Map operations is replaced with string concatenation, multiplication in Map is replaced with character replication, zero-padding in Pad becomes character deletion, and the numbers 00-$9$ are replaced with the characters $a$-$j$.
For humans, although the process of reasoning tends to be implicit, according to the traditional definition in logic, it can be divided into two stages: induction and deduction. Here, we focus on the objectives of these two stages: deriving the correct rules through induction and making inferences on new instances using deduction. We aim to explore the correlation between the two during the reasoning process of LLMs. Therefore, for each scenario, we design two tasks: rule induction (RI) and example inference (EI), defined as follows:
  • Rule Induction Task: Given an observed fact set ${\mathbb{X}}$, this task evaluates the model’s accuracy in inducing transformation rules $f$. It aims to evaluate the model’s proficiency in mastering intermediate rules during inductive reasoning (Yang et al., 2024b; Shao et al., 2024).
  • Example Inference Task: Given an observed fact set ${\mathbb{X}}$, the task provides an unseen example input ${\bm{x}}_{t}$ as input and measures the accuracy of the predicted ${\bm{y}}_{t}$. It aims to evaluate the final performance of the model’s inductive reasoning (Chollet, 2019; Xu et al., 2024a; Qiu et al., 2024).
We provide all the prompts used for these tasks in Appendix A.3.
@@ 3.1 Overall Performances on Mirage
Llama2-13B RI 0.01 0.00 0.00 0.03 0.01 0.01 0.00 0.21 0.00 0.01 0.00 0.10
EI 0.26 0.11 0.25 0.22 0.13 0.03 0.14 0.25 0.06 0.01 0.06 0.19
Llama3-8B RI 0.15 0.11 0.19 0.19 0.23 0.04 0.14 0.22 0.16 0.02 0.08 0.21
EI 0.30 0.15 0.25 0.25 0.20 0.12 0.25 0.29 0.09 0.11 0.16 0.24
GPT-4o RI 0.41 0.32 0.38 0.32 0.35 0.21 0.44 0.30 0.33 0.16 0.41 0.24
EI 0.68 0.37 0.61 0.56 0.58 0.25 0.64 0.39 0.42 0.17 0.49 0.29
GPT-4 RI 0.47 0.29 0.41 0.28 0.58 0.22 0.56 0.27 0.46 0.15 0.45 0.23
EI 0.68 0.37 0.61 0.57 0.63 0.29 0.71 0.44 0.42 0.21 0.64 0.30
Claude-3.5 RI 0.44 0.35 0.34 0.46 0.22 0.20 0.38 0.33 0.24 0.13 0.38 0.26
EI 0.79 0.45 0.62 0.58 0.65 0.33 0.76 0.45 0.46 0.24 0.59 0.30
Table 1: Overall performance (accuracy) of different models on Mirage. The best results in rule induction (RI) are in bold, while the best results in example inference (EI) are underlined.
@@ Results
The results are shown in Table 1, from which we can draw the following conclusions: (1) LLMs’ inductive reasoning does not rely on rule induction. Given the same set of observed facts, the model’s performance on rule induction is noticeably worse than on example inference in almost all cases. This suggests that most of the model’s correct deductions do not depend on inducing a correct rule. (2) LLMs face difficulties in handling inductive reasoning in real-world problems. When comparing different scenarios, all models perform the worst on the RP tasks. For example, GPT-4o only achieves 0.16 and 0.17 accuracy when the dimension is 8. This indicates that, compared to purely symbolic forms (LT, CG, ST), natural language forms pose a greater challenge for the models’ inductive reasoning abilities.
@@
In the main experiments, we find that there is a significant performance gap between rule induction and example inference for LLMs. However, this gap may be caused by the difference in difficulty between the two tasks. When the model is unable to perform correct inductive reasoning,
Model Rule Induction Example Inference
@@
it is likely to guess the correct answers for the EI task more easily compared to the RI task, resulting in a higher accuracy. We conduct this additional experiment to eliminate the interference of this factor. Specifically, we randomly perturb one fact in ${\mathbb{X}}$ to violate rule $f$. Then, we observe the performance of both tasks and calculate the change rate (CR) of accuracy before and after the perturbation. CR represents the sensitivity of the model’s performance to the input. If CR is high, it indicates a strong correlation between task performance and input, making it difficult to answer the question correctly through random guessing, Therefore, CR can serve as an indicator of the reasoning difficulty for the task. We randomly choose 100 pieces of test data from the dataset and generate questions under the LF scenario. [^4] Unless otherwise specified, this configuration will be maintained for all subsequent experiments. The experimental results on different models are demonstrated in Table 2. We can observe that the two tasks have comparable CR for both models, indicating that the reasoning difficulty of the EI task is not lower than the RI task. The tasks themselves do not cause such a large performance gap.
@@ 3.2 Performances of Advanced Methods
In §$\S$3.1, we observe that LLMs perform poorly on our dataset, especially in rule induction tasks. Considering previous work has proposed numerous methods to elicit the model’s reasoning abilities (Wei et al., 2022; Wang et al., 2023b; Madaan et al., 2023), we wonder whether they can boost models’ performance on Mirage.
@@
Since we focus on exploring the model’s intrinsic capabilities, we only consider methods that do not introduce any external tools or knowledge. Specifically, the methods are as follows: Input-Output (IO): We prompt models to generate answers directly under different shots. Inductive-Deductive (ID): We prompt models to generate rules for RI and apply them to answer questions in EI. Chain-of-Thought (CoT) (Wei et al., 2022): We prompt models to generate rationales and answers for the two tasks. Self-Consistency (SC) (Wang et al., 2023b): Based on CoT, we sample $n$ rationales and use the major voting strategy to predict the final answer. Self-Refine (SR) (Madaan et al., 2023): We prompt the model to provide feedback on the generated rules, and then refine the rules based on that feedback (with a maximum of $t$ iterations). After the iteration stops, we use the latest rule to answer the RI and apply it to answer the EI. Hypothesis Refinement (HR) (Qiu et al., 2024): HR is an optimized version of SR, which first generates $n$ rules. In each iteration, we apply the current rules to all observed examples, compare the actual output with the expected output, and get the number of correct predictions along with the information about incorrect examples. If a candidate rule is correct for all observed facts, it is immediately returned. Otherwise, the rule with the highest number of correct predictions and the associated error information is used as input for the model to refine, generating $n$ rules for the next round, until the maximum number of iterations $t$ is reached. We sample 200 questions for each test.
@@
RI EI ($\Delta$) RI EI ($\Delta$) RI EI ($\Delta$) RI EI ($\Delta$)